OpenStudy (anonymous):

Hi everyone!! Happy New Year to you all! :)can somebody PLEASE help me?!! how do you simplify 16^1/2?? how do you describe in a paragraph how you simplify the squareroot of 100(i know the answer is 10 nbut how do you desribe how you got the answer?)

OpenStudy (anonymous):

16^1=16 16/2=8/1=8

OpenStudy (anonymous):

if you meant 16 to the power of a half, the answer is 4, becasue a power of a half is the same as the sqrt

OpenStudy (anonymous):

hi Zara the question is how do you simplify 16 with the squareroot of one half not 1 & 2 but thank you for responding

OpenStudy (anonymous):

ohh, sorry :)

OpenStudy (anonymous):

thank you 08krishna,o you know how i would describe how to get the answer in a paragraph?

OpenStudy (anonymous):

zara no need to be sorry, thank you for responding

OpenStudy (anonymous):

haha thanks, no problem :)

OpenStudy (anonymous):

\[\Large {x^{n/m}} = \sqrt[m]{{{x^n}}}\]